# Security modelling for cyber-physical systems: A systematic literature review

### v1

@@ 6.2 Future work
We may also use Petri net based approaches (Chen et al. (2011); Dahl and Wolthusen (2006)) to better model CPS intrusions, integrating both cyber and physical actions. Petri nets are a graphical and mathematical tool for studying concurrent and distributed systems and have been widely applied in many different areas of computer science and other disciplines (Murata (1989)). In the context of CPS intrusion analysis, Petri nets are invaluable for modelling the interactions between cyber and physical components over time, thereby overcoming limitations of existing security modelling approaches when considering dynamic, multi-agent cyber-physical threats.
@@ 6.3 Limitations of this review
Finally, our research questions were focused on how existing security modelling approaches addressed CPS, and corresponding adversarial tactics, techniques, and procedures. These factors may introduce some bias to the articles selected for review and may miss out articles that are relevant to CPS but did not mention adversarial models specifically. We addressed this by reviewing bibliographies in the selected articles and then including additional articles that may be relevant to the review.
@@ 7 Conclusion
In this review, we identified a number of state-of-the-art threat modelling and attack modelling studies relevant to cyber-physical systems (CPS), with the following findings.
  • Threat modelling is commonly conducted in early stages of system development. However, this means that evolving attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles will render threat models outdated and irrelevant.
  • Security models in the literature generally focus on IT systems, and these models are challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
  • Most papers do not differentiate between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks, effects and consequences are considered in security modelling.
  • There is ambiguity in the literature regarding the definitions and relationship between threat modelling and attack modelling. Correspondingly, a unified security modelling framework that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article.

### v2

@@ 6.2 Future work
We may also use Petri net based approaches [55] [56] to model CPS intrusions better by integrating both cyber and physical actions. Petri nets are a graphical and mathematical tool for studying concurrent and distributed systems and have been widely applied in many different areas of computer science and other disciplines [57]. In the context of CPS intrusion analysis, Petri nets are invaluable for modelling the interactions between cyber and physical components over time, thereby overcoming limitations of existing security modelling approaches when considering dynamic, multi-agent cyber-physical threats.
@@ 6.3 Limitations of this review
Our research questions were focused on how existing security modelling approaches addressed CPS, and corresponding adversarial tactics, techniques, and procedures. These factors may introduce some bias to the articles selected for review and may miss out articles that are relevant to CPS but did not mention adversarial models specifically. We addressed this by reviewing bibliographies in the selected articles and then including additional articles that may be relevant to the review.
@@ 7 Conclusion
In this review, we identified several state-of-the-art threat modelling and attack modelling studies relevant to CPS, with the following findings:
  • Threat modelling is usually conducted in early stages of system development, and unlikely to anticipate new attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles.
  • Security models in the literature that focus on IT systems could be challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
  • There is limited differentiation between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks and consequences are considered in security modelling.
  • Unclear definitions and relationships between threat modelling and attack modelling can lead to inconsistent security modelling approaches. A unified security modelling approach that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article to address this.
@@
Declarations
The authors make the following declarations about this work:
Availability of data and materials. This article does not involve the generation of any original datasets or materials. All data and resources referenced in this study are publicly available and can be accessed through the sources cited within the article.
Competing interests. The authors have no competing interests to declare.
Funding. The authors did not receive support from any organisation for the submitted work.
Authors’ contributions. All authors contributed to the study conception and design. The literature review and analysis were performed by Shaofei Huang. The first draft of the manuscript was prepared by Shaofei Huang and all authors reviewed and edited previous versions of the manuscript. All authors read and approved the final manuscript.
Acknowledgements. The authors have no acknowledgments to declare.